# Numerically “exact” charge transport dynamics in a dissipative electron–phonon model rationalizing the success of the transient localization scenario

### v1

@@ IV.1 Influence of damping on transport dynamics
Figure 5: Dynamics of (a) the diffusion constant $\mathcal{D}$ and (b) the diffusion exponent $\alpha$ for different values of the friction coefficient $\gamma_{0}$. (c) The dynamical-mobility profile $\mathrm{Re}\>\mu(\omega)$ for different values of $\gamma_{0}$. The TLS predictions employ $\alpha_{d}=2.2$. The insets in (a) and (b) display the dynamics of $\mathcal{D}$ and $\alpha$, respectively, on intermediate-to-long timescales. The model parameters are $J=1,\omega_{0}=0.044,\lambda=0.336,T=0.175$. DEOM computations are performed for $N=31,D=4$. Note the logarithmic scale on horizontal axes.
In Figs. 5(a)–5(c), we monitor the changes to transport quantities brought about by the damping of vibrational motions.
Figures 5(a) and 5(b) reveal that a realistically strong damping of vibrational motions starts to influence carrier dynamics on intermediate timescales (for $t\gtrsim t_{\mathrm{min}}\approx 1/\omega_{0}$). Therefore, realistically strong damping does not change the position and weakly affects the shape of the displaced Drude peak, see Fig. 5(c), which reflects short-time carrier dynamics in the field of effectively frozen vibrations. [91, 3, 15] The timescale $\tau_{\mathrm{diff}}$ on which the diffusive carrier dynamics is established can be estimated as the characteristic timescale for the decay of the current autocorrelation function, which can be defined as [see also Eq. (47)]
@@
Equation (43) involves the long-time diffusion constant, which somewhat decreases with damping, [68] see the inset of Fig. 5(a), and the thermodynamic quantity $\langle j^{2}\rangle$, which is expected to be virtually independent of the damping, see Sec. III.2. While the approach towards diffusive transport becomes somewhat more rapid with increasing damping, the order of magnitude of $\tau_{\mathrm{diff}}$ remains constant for realistic values of $\gamma_{0}$.
Increasing $\gamma_{0}$ renders the intermediate-time growth of $\mathcal{D}(t)$ less pronounced until it is fully suppressed for $\gamma_{0}/\omega_{0}\gtrsim 0.5$, see the inset of Fig. 5(a). Interestingly, the magnitude of the growth, quantified by $\mathcal{D}_{\infty}-\mathcal{D}(t_{\mathrm{min}})$, determines whether the super-to-subdiffusive crossover is transient or permanent, see Fig. 5(b), as well as the low-frequency features of the dynamical mobility, see Fig. 5(c). For $\gamma_{0}/\omega_{0}=0.05$ and 0.15, $\mathcal{D}_{\infty}-\mathcal{D}(t_{\mathrm{min}})$ is sufficiently large so that the diffusive transport sets in from the superdiffusive side, giving rise to the enhancement in the dynamical mobility below the vibrational frequency. Although $\mathcal{D}_{\infty}-\mathcal{D}(t_{\mathrm{min}})$ remains positive for $\gamma_{0}/\omega_{0}=0.25$, we find that its magnitude is such that the diffusive transport is approached from the subdiffusive side, while there is almost no low-frequency enhancement in the dynamical mobility, see also Fig. 6(b).
@@ IV.2 Success of the transient localization phenomenology
We have used that the expectation value of the operator $g_{\xi}(b_{n\xi}^{\dagger}+b_{n\xi})$ is distributed according to the Gaussian distribution of zero mean and variance $g_{\xi}^{2}\coth(\beta\omega_{\xi}/2)\approx 2g_{\xi}^{2}/(\beta\omega_{\xi})$[91] The approximate equality holds for $\beta\omega_{\xi}\ll 1$, as appropriate in the high-temperature slow-phonon regime (see Sec. II.6) and on short timescales we consider. As the operators $g_{\xi}(b_{n\xi}^{\dagger}+b_{n\xi})$ commute for different values of $n$ or $\xi$, the expectation value of their sum over $\xi$ is distributed in the same way as the sum of mutually independent Gaussian variables, each of which has zero mean and variance $2g_{\xi}^{2}/(\beta\omega_{\xi})$, thus justifying Eq. (46). Equation (46) additionally reveals that the reference statically disordered model does not depend on the details of the carrier–phonon SD or phonon dynamics [Eq. (11)], but only on the overall interaction strength quantified by the dimensionless parameter $\lambda$. This once again reflects the fact that the statically disordered model is appropriate only on short timescales.
@@
The timescale $\tau_{d}$ in Eq. (44), after which phonon dynamics cannot be neglected, is assumed as $\tau_{d}^{-1}=\alpha_{d}\omega_{0}$[15, 3] where the free parameter $\alpha_{d}\sim 1$ can be finetuned to best reproduce some reference (experimental or numerical) result for $\mu_{\mathrm{dc}}$ without qualitatively affecting TLS predictions for transport dynamics. In Figs. 5(a)–5(c), we show the TLS result using $\alpha_{d}=2.2$. This value was suggested by comparing TLS and reference numerically exact results in Refs. 95, 96, which considered undamped vibrations. Here, this standard choice for $\alpha_{d}$ gives $\mu_{\mathrm{dc}}^{\mathrm{TLS}}$ that differs by at most $20\%$ from the corresponding DEOM results for physically reasonable damping strengths, see the inset of Fig. 5(a). Importantly, both the TLS with $\alpha_{d}=2.2$ and DEOM with physically relevant values of $\gamma_{0}$ yield $\mu_{\mathrm{dc}}$ that falls into the experimental range $\mu_{\mathrm{dc}}^{\mathrm{exp}}\sim(10-20)\>\mathrm{cm}^{2}/(\mathrm{Vs})$ for the room-temperature dc mobility in rubrene [97, 98, 99] [our DEOM mobility is measured in units of $e_{0}a_{l}^{2}/\hbar$, corresponding to $7.9\>\mathrm{cm}^{2}/(\mathrm{Vs})$ for $a_{l}=7.2\mathrm{\AA }$, see, e.g., Ref. 23].
@@
Therefore, Figs. 5(a)–5(c) strongly suggest that the damping in intermolecular motions, which is inevitably present in a realistic material, [23, 30, 31, 32, 76] is behind the success of the transient-localization phenomenology (and the related Drude–Anderson framework) in explaining the experimental carriers’ optical response. Notably, the damping does not have to be excessively strong in order for DEOM results to comply with the transient-localization phenomenology. Figures 5(a)–5(c) reveal that this happens already for $\gamma_{0}/\omega_{0}\gtrsim 0.25$, which belongs to the underdamped regime characterized by a pronounced peak in the SD, see Fig. 1. We thus propose that the optical-absorption enhancements below the vibrational frequency reported in model-Hamiltonian studies assuming undamped vibrations [25, 90, 91, 21, 92, 93] are to be regarded as artifacts of undamped vibrations.
@@ IV.3 A closer look into the TLS phenomenological timescale
Here, we connect the parameter $\tau_{d}$ in Eq. (44) with the characteristic timescale $\tau_{c}$ of phonon-induced fluctuations of the transfer integral. In the dissipative model studied here, $\tau_{c}$ can be estimated as [60]
@@
where $\Psi(t)=-2\int_{t}^{+\infty}ds\>\mathrm{Im}\>\mathcal{C}(s)$ is the so-called relaxation function. [48, 60] For $\gamma_{0}/\omega_{0}\gtrsim 0.25$, for which $\alpha(t)$ and $\mathrm{Re}\>\mu(\omega)$ display TL phenomenology, we extract $\alpha_{d}=(\omega_{0}\tau_{d})^{-1}$ by requiring that the dc mobility within the TLS is identical to the corresponding DEOM result. Figure 6(a) shows that thus extracted $\alpha_{d}$ is correlated with $\tau_{c}$, see the powerlaw fit (solid line) whose parameters are summarized in the caption. Interestingly, $\tau_{d}=(\alpha_{d}\omega_{0})^{-1}$ somewhat increases with the damping, without changing its order of magnitude, contrary to the timescale $\tau_{\mathrm{diff}}$ on which carrier dynamics become diffusive, see Sec. IV.1. [Since we choose $\tau_{d}$ so that $\mu_{\mathrm{dc}}^{\mathrm{TLS}}=\mu_{\mathrm{dc}}^{\mathrm{DEOM}}$, and since $C_{jj}(0)$ is identical within TLS and DEOM, Eq. (43) shows that $\tau_{\mathrm{diff}}$ is also identical within TLS and DEOM.] Such a trend in $\tau_{d}$ with increasing $\gamma_{0}$ is compatible with the dc mobility decreasing with $\gamma_{0}$, see Figs. 5(a) and 5(c). Namely, as vibrations become more damped, the model becomes increasingly similar to that with only static disorder, [31] and this reference point of the TLS is gradually approached as $\tau_{d}$ is increased, see Eq. (44). This is discussed in more detail in Appendix B.
If we choose $\alpha_{d}$ so that the TLS perfectly reproduces DEOM mobility, the overall agreement between TLS and DEOM dynamical-mobility profiles improves with increasing $\gamma_{0}$, see Fig. 6(b). At smaller damping strengths, the requirement $\mu_{\mathrm{dc}}^{\mathrm{TLS}}=\mu_{\mathrm{dc}}^{\mathrm{DEOM}}$ makes the TLS almost completely miss the DDP, see the curves for $\gamma_{0}/\omega_{0}=0.25$ in Fig. 6(b). Although a more accurate TLS approximation to the numerically exact DDP would underestimate the dc mobility, $\mu_{\mathrm{dc}}^{\mathrm{TLS}}$ would still be of the same order of magnitude as $\mu_{\mathrm{dc}}^{\mathrm{DEOM}}$. Meanwhile, at larger damping strengths, the TLS can reproduce reasonably well both $\mu_{\mathrm{dc}}$ and the position, intensity, and shape of the DDP, see the curves for $\gamma_{0}/\omega_{0}=1.7$ in Fig. 6(b). For even stronger dampings, we expect that the dependence of $\tau_{d}$ on $\gamma_{0}$ is different from that in Fig. 6(a). In Appendix B, we argue that $\tau_{d}$ should depend linearly on $\gamma_{0}$ in the strong-damping limit.
@@ IV.4 Transport dynamics for the Drude–Lorentz spectral density
Considering the DL SD, the model parameters are usually chosen as in Ref. 40, so that $J=300\>\mathrm{cm}^{-1}$, $2E_{0}=323\>\mathrm{cm}^{-1}$, and $\gamma_{\mathrm{DL}}=41\>\mathrm{cm}^{-1}$. However, the models considered in Refs. 40, 44 somewhat differ from the model we consider. Namely, the nearest-neighbor hopping amplitude in Ref. 40 (Ref. 44) is modulated by the displacement of the oscillator on the left end of a bond (the center of a bond), so that the dimensionless interaction constant is $\lambda=\frac{E_{0}}{J}$. This is in contrast to our definition $\lambda=\frac{2E_{0}}{J}$, which reflects the fact that the hopping amplitude is modulated by the difference of the displacements on two ends of a bond. Importantly, $\lambda\approx 0.5$ in Refs. 40, 44, which belongs to the intermediate-interaction regime relevant for organic molecular crystals. [5]
Figures 7(a) and 7(b) present DEOM results for Ishizaki’s BO approximation to the DL SD [60] assuming that $J=300\>\mathrm{cm}^{-1}$, $E_{0}=75\>\mathrm{cm}^{-1}$ ($\lambda=0.5$), $\gamma_{0}=2\gamma_{\mathrm{DL}}=80\>\mathrm{cm}^{-1}$ ($\gamma_{0}/J=4/15$), $\epsilon=\gamma_{0}/4=20\>\mathrm{cm}^{-1}$ ($\gamma_{0}/\omega_{0}=4/\sqrt{17}$), at room temperature $T=210\>\mathrm{cm}^{-1}$ ($T/J=0.7$). DEOM computations are performed using $N=11$, $D=8$, and $K=2$. We have checked that increasing $D$ from 8 to 9 or retaining more than two terms in the exponential decomposition of Eq. (11) ($K=3$ or 4 with $N=7$) introduces only minor quantitative changes to the DEOM result for $N=11,D=8,K=2$. We have also checked that employing a smaller value of $\epsilon=\gamma_{0}/20$ does not change (either qualitatively or quantitatively) the results using $\epsilon=\gamma_{0}/4$.
@@
Finally, for the usual choice of model parameters of the DL SD at room temperature, [40, 44] the phonon-assisted contribution to $\mu_{\mathrm{dc}}$ [obtained by considering only $\langle j_{\mathrm{e-ph}}(t)j_{\mathrm{e-ph}}(0)\rangle$ in Eq. (32)] is comparable to the purely electronic contribution to $\mu_{\mathrm{dc}}$ [obtained by considering only $\langle j_{\mathrm{e}}(t)j_{\mathrm{e}}(0)\rangle$ in Eq. (32)], see Fig. 7(b). This is different from the situation for model parameters we use with BO SD at room temperature, for which the phonon-assisted contribution is much smaller than the purely electronic contribution to $\mu_{\mathrm{dc}}$ [25] irrespective of the damping analyzed. The reason behind this difference lies in the different values of the transfer integral considered in the two models. Namely, the authors of Ref. 40 suggest that the Holstein-type interaction with high-frequency intramolecular vibrations leads to an effective reduction of the hopping amplitude from its bare value $J\approx 1150\>\mathrm{cm}^{-1}$ (see, e.g.,  23) to the dressed value $J\approx 300\>\mathrm{cm}^{-1}$. Keeping this in mind, and remembering that the value of $\lambda$ used with the DL SD (0.5) is somewhat larger than that used with the BO SD (0.336), it is not surprising that the dc mobility in the parameter regime discussed in Refs. 40, 44 is almost an order of magnitude smaller than in the regime analyzed in Fig. 5, compare Fig. 7(b) to Fig. 5(c).

### v2

@@ IV.1 Influence of damping on transport dynamics
Figure 5: Dynamics of (a) the diffusion constant $\mathcal{D}$ and (b) the diffusion exponent $\alpha$ for different values of the friction coefficient $\gamma_{0}$. (c) The dynamical-mobility profile $\mathrm{Re}\>\mu(\omega)$ for different values of $\gamma_{0}$. The TLS predictions employ $\alpha_{d}=2.2$. The insets in (a) and (b) display the dynamics of $\mathcal{D}$ and $\alpha$, respectively, on intermediate-to-long timescales. The model parameters are $J=1,\omega_{0}=0.044,\lambda=0.336,$ and $T=0.175$. DEOM computations are performed for $N=31,D=4$. Note the logarithmic scale on the horizontal axes.
In Figs. 5(a)–5(c), we monitor the changes to transport quantities brought about by the nonzero width of the spectrum of relevant phonons, which is embodied in the damping parameter $\gamma_{0}$ of the present model [Eq. (6)].
Figures 5(a) and 5(b) reveal that a realistically strong damping of vibrational motions starts to influence carrier dynamics on intermediate timescales (for $t\gtrsim t_{\mathrm{min}}\approx 1/\omega_{0}$). Therefore, realistically strong damping (a realistically wide phonon spectrum) does not change the position and weakly affects the shape of the displaced Drude peak, which reflects short-time carrier dynamics in the field of effectively frozen vibrations [91, 3, 15, 94] [see Fig. 5(c)]. The timescale $\tau_{\mathrm{diff}}$ on which the diffusive carrier dynamics is established can be estimated as the characteristic timescale for the decay of the current autocorrelation function, which can be defined as [see also Eq. (47)]
@@
Equation (43) involves the long-time diffusion constant, which somewhat decreases with damping [87] [see the inset of Fig. 5(a)], and the thermodynamic quantity $\langle j^{2}\rangle$, which is expected to be virtually independent of the damping (see Sec. III.2). While the approach toward diffusive transport becomes somewhat more rapid with increasing damping, the order of magnitude of $\tau_{\mathrm{diff}}$ remains constant for realistic values of $\gamma_{0}$.
Increasing $\gamma_{0}$ renders the intermediate-time growth of $\mathcal{D}(t)$ less pronounced until it is fully suppressed for $\gamma_{0}/\omega_{0}\gtrsim 0.5$ [see the inset of Fig. 5(a)]. Interestingly, the magnitude of the growth, quantified by $\mathcal{D}_{\infty}-\mathcal{D}(t_{\mathrm{min}})$, determines whether the super-to-subdiffusive crossover is transient or permanent [see Fig. 5(b)] as well as the low-frequency features of the dynamical mobility [see Fig. 5(c)]. For $\gamma_{0}/\omega_{0}=0.05$ and 0.15, $\mathcal{D}_{\infty}-\mathcal{D}(t_{\mathrm{min}})$ is sufficiently large so that the diffusive transport sets in from the superdiffusive side, giving rise to the enhancement in the dynamical mobility below the vibrational frequency. Although $\mathcal{D}_{\infty}-\mathcal{D}(t_{\mathrm{min}})$ remains positive for $\gamma_{0}/\omega_{0}=0.25$, we find that its magnitude is such that the diffusive transport is approached from the subdiffusive side, while there is almost no low-frequency enhancement in the dynamical mobility [see also Fig. 6(b)].
@@ IV.2 Success of the transient localization phenomenology
We have used that the expectation value of the operator $g_{\xi}(b_{n\xi}^{\dagger}+b_{n\xi})$ is distributed according to the Gaussian distribution of zero mean and variance $g_{\xi}^{2}\coth(\beta\omega_{\xi}/2)\approx 2g_{\xi}^{2}/(\beta\omega_{\xi})$[91] The approximate equality holds for $\beta\omega_{\xi}\ll 1$, as appropriate in the high-temperature slow-phonon regime (see Sec. II.6) and on short timescales we consider. As the operators $g_{\xi}(b_{n\xi}^{\dagger}+b_{n\xi})$ commute for different values of $n$ or $\xi$, the expectation value of their sum over $\xi$ is distributed in the same way as the sum of mutually independent Gaussian variables, each of which has zero mean and variance $2g_{\xi}^{2}/(\beta\omega_{\xi})$, thus justifying Eq. (46). Equation (46) additionally reveals that the reference statically disordered model does not depend on the details of the carrier–phonon SD or phonon dynamics [Eq. (11)] but only on the overall interaction strength quantified by the dimensionless parameter $\lambda$ [Eq. (8)]. This once again reflects the fact that the statically disordered model is appropriate only on short timescales.
@@
In Eq. (44), the timescale $\tau_{d}$, after which phonon dynamics cannot be neglected, is assumed as $\tau_{d}^{-1}=\alpha_{d}\omega_{0}$[15, 3] where the free parameter $\alpha_{d}\sim 1$ can be finetuned to best reproduce some reference (experimental or numerical) result for $\mu_{\mathrm{dc}}$ without qualitatively affecting TLS predictions for transport dynamics. In Figs. 5(a)–5(c), we show the TLS result using $\alpha_{d}=2.2$. This value was suggested by comparing TLS and reference numerically exact results in Refs. 65, 66, which considered undamped vibrations. Here, this standard choice for $\alpha_{d}$ gives $\mu_{\mathrm{dc}}^{\mathrm{TLS}}$ that differs by at most $20\%$ from the corresponding DEOM results for physically reasonable damping strengths [see the inset of Fig. 5(a)]. Importantly, both the TLS with $\alpha_{d}=2.2$ and DEOM with physically relevant values of $\gamma_{0}$ yield $\mu_{\mathrm{dc}}$ that falls into the experimental range $\mu_{\mathrm{dc}}^{\mathrm{exp}}\sim(10-20)\>\mathrm{cm}^{2}/(\mathrm{Vs})$ for the room-temperature dc mobility in rubrene [96, 97, 98] [our DEOM mobility is measured in units of $e_{0}a_{l}^{2}/\hbar$, corresponding to $7.9\>\mathrm{cm}^{2}/(\mathrm{Vs})$ for $a_{l}=7.2\mathrm{\AA }$, see, e.g., Ref. 23].
@@
Therefore, Figs. 5(a)–5(c) strongly suggest that a nonzero width of the spectrum of relevant phonons, [23, 22, 32] which in the present model amounts to a nonzero damping in intermolecular motions, [30, 31, 71] is behind the success of the transient-localization phenomenology (and the related Drude–Anderson framework) in explaining the experimental carriers’ optical response. Notably, the damping does not have to be excessively strong (the relevant-phonon spectrum does not have to be excessively wide) in order for DEOM results to comply with the transient-localization phenomenology. Figures 5(a)–5(c) reveal that this happens already for $\gamma_{0}/\omega_{0}\gtrsim 0.25$, which belongs to the underdamped regime characterized by a pronounced peak in the SD (see Fig. 1). We thus propose that the optical-absorption enhancements below the vibrational frequency reported in model-Hamiltonian studies [25, 90, 91, 21, 92, 93] are to be regarded as artifacts of the delta-like phonon spectrum (undamped vibrations).
@@ IV.3 A closer look into the TLS phenomenological timescale
The parameter $\tau_{d}$ in Eq. (44) is associated with the characteristic timescale $\tau_{c}$ of phonon-induced fluctuations of the transfer integral. In models with undamped phonons and in the limit of zero carrier density, it exhibits undamped oscillations with phonon frequency $\omega_{0}$ [99, 88] so that $\tau_{c}\propto\omega_{0}^{-1}$. However, the fact that Eq. (44) features an exponentially decaying (and not an oscillatory) factor suggests that the TLS is even more appropriate for dissipative models, in which the lifetime of transfer-integral fluctuations is finite. In our model, $\tau_{c}$ can be estimated as [62]
@@
where $\Psi(t)=-2\int_{t}^{+\infty}ds\>\mathrm{Im}\>\mathcal{C}(s)$ is the so-called relaxation function. [48, 62]
For $\gamma_{0}/\omega_{0}\gtrsim 0.25$, for which $\alpha(t)$ and $\mathrm{Re}\>\mu(\omega)$ display TL phenomenology, we extract $\alpha_{d}=(\omega_{0}\tau_{d})^{-1}$ by requiring that the dc mobility within the TLS is identical to the corresponding DEOM result. Figure 6(a) shows that the extracted $\alpha_{d}$ is correlated with $\tau_{c}$, see the powerlaw fit (solid line) whose parameters are summarized in the caption. Interestingly, $\tau_{d}=(\alpha_{d}\omega_{0})^{-1}$ somewhat increases with the damping, without changing its order of magnitude, contrary to the timescale $\tau_{\mathrm{diff}}$ on which carrier dynamics become diffusive (see Sec. IV.1). [Since we choose $\tau_{d}$ so that $\mu_{\mathrm{dc}}^{\mathrm{TLS}}=\mu_{\mathrm{dc}}^{\mathrm{DEOM}}$, and since $C_{jj}(0)$ is identical within TLS and DEOM, Eq. (43) shows that $\tau_{\mathrm{diff}}$ is also identical within TLS and DEOM.] Such a trend in $\tau_{d}$ with increasing $\gamma_{0}$ is compatible with the dc mobility decreasing with $\gamma_{0}$, see Figs. 5(a) and 5(c). In particular, as vibrations become more damped, the model becomes increasingly similar to that with only static disorder, [31] and this reference point of the TLS is gradually approached as $\tau_{d}$ is increased [see Eq. (44)]. This is discussed in more detail in Appendix B.
If we choose $\alpha_{d}$ so that the TLS perfectly reproduces DEOM mobility, the overall agreement between TLS and DEOM dynamical-mobility profiles improves with increasing $\gamma_{0}$ [see Fig. 6(b)]. At smaller damping strengths, the requirement $\mu_{\mathrm{dc}}^{\mathrm{TLS}}=\mu_{\mathrm{dc}}^{\mathrm{DEOM}}$ makes the TLS almost completely miss the DDP [see the curves for $\gamma_{0}/\omega_{0}=0.25$ in Fig. 6(b)]. Although a more accurate TLS approximation to the numerically exact DDP would underestimate the dc mobility, $\mu_{\mathrm{dc}}^{\mathrm{TLS}}$ would still be of the same order of magnitude as $\mu_{\mathrm{dc}}^{\mathrm{DEOM}}$. Meanwhile, at larger damping strengths, the TLS can reproduce reasonably well both $\mu_{\mathrm{dc}}$ and the position, intensity, and shape of the DDP [see the curves for $\gamma_{0}/\omega_{0}=1.7$ in Fig. 6(b)]. For even stronger dampings, we expect that the dependence of $\tau_{d}$ on $\gamma_{0}$ is different from that in Fig. 6(a). In Appendix B, we argue that $\tau_{d}$ should depend linearly on $\gamma_{0}$ in the strong-damping limit.
@@ IV.4 Transport dynamics for the Drude–Lorentz spectral density
Considering the DL SD, the model parameters are usually chosen as in Ref. 40, so that $J=300\>\mathrm{cm}^{-1}$, $2E_{0}=323\>\mathrm{cm}^{-1}$, and $\gamma_{\mathrm{DL}}=41\>\mathrm{cm}^{-1}$. However, the models considered in Refs. 40, 44 somewhat differ from the model we consider. In particular, the nearest-neighbor hopping amplitude in Ref. 40 (Ref. 44) is modulated by the displacement of the oscillator on the left end of a bond (the center of a bond) so that the dimensionless interaction constant is $\lambda=\frac{E_{0}}{J}$. This is in contrast to our definition $\lambda=\frac{2E_{0}}{J}$, which reflects the fact that the hopping amplitude is modulated by the difference of the displacements on two ends of a bond. Importantly, $\lambda\approx 0.5$ in Refs. 40, 44, which belongs to the intermediate-interaction regime relevant for organic molecular crystals. [5]
Figures 7(a) and 7(b) present DEOM results for Ishizaki’s BO approximation to the DL SD [62] assuming that $J=300\>\mathrm{cm}^{-1}$, $E_{0}=75\>\mathrm{cm}^{-1}$ ($\lambda=0.5$), $\gamma_{0}=2\gamma_{\mathrm{DL}}=80\>\mathrm{cm}^{-1}$ ($\gamma_{0}/J=4/15$), and $\epsilon=\gamma_{0}/4=20\>\mathrm{cm}^{-1}$ ($\gamma_{0}/\omega_{0}=4/\sqrt{17}$), at room temperature $T=210\>\mathrm{cm}^{-1}$ ($T/J=0.7$). DEOM computations are performed using $N=11$, $D=8$, and $K=2$. We have checked that increasing $D$ from 8 to 9 or retaining more than two terms in the exponential decomposition of Eq. (11) ($K=3$ or 4 with $N=7$) introduces only minor quantitative changes to the DEOM result for $N=11,D=8,K=2$. We have also checked that employing a smaller value of $\epsilon=\gamma_{0}/20$ does not change (either qualitatively or quantitatively) the results using $\epsilon=\gamma_{0}/4$.
@@
Finally, for the usual choice of model parameters of the DL SD at room temperature, [40, 44] the phonon-assisted contribution to $\mu_{\mathrm{dc}}$ [obtained by considering only $\langle j_{\mathrm{e-ph}}(t)j_{\mathrm{e-ph}}(0)\rangle$ in Eq. (32)] is comparable to the purely electronic contribution to $\mu_{\mathrm{dc}}$ [obtained by considering only $\langle j_{\mathrm{e}}(t)j_{\mathrm{e}}(0)\rangle$ in Eq. (32)] [see Fig. 7(b)]. This is different from the situation for model parameters we use with BO SD at room temperature, for which the phonon-assisted contribution is much smaller than the purely electronic contribution to $\mu_{\mathrm{dc}}$[25] irrespective of the damping analyzed. The reason behind this difference lies in the different values of the transfer integral considered in the two models. In particular, the authors of Ref. 40 suggest that the Holstein-type interaction with high-frequency intramolecular vibrations leads to an effective reduction of the hopping amplitude from its bare value $J\approx 1150\>\mathrm{cm}^{-1}$ (see, e.g., Ref. 23) to the dressed value $J\approx 300\>\mathrm{cm}^{-1}$. Keeping this in mind and remembering that the value of $\lambda$ used with the DL SD (0.5) is somewhat larger than that used with the BO SD (0.336), it is not surprising that the dc mobility in the parameter regime discussed in Refs. 40, 44 is almost an order of magnitude smaller than in the regime analyzed in Fig. 5 [compare Fig. 7(b) to Fig. 5(c)].